# Flavor Patterns of Fundamental Particles from Quantum Entanglement?

### v1

@@ Appendix A Description of Scattering/Measurement Process
The $\mathcal{S}$ operator is unitary and it can defined in terms of the transfer operator $\mathcal{T}$:
@@
    $\Pi_{\rm{out}}$ corresponds to the final state measurement. First, we restrict the final state to have two particles [14] with negative helicity, a combination of actions denoted by $\Pi_{2}^{L}$. Then, we measure the scattering angle $\Theta$ of one of the final state particles, which without loss of generality we choose to be the up-type quark. The full measurement can be written schematically as:
    $\Pi_{\rm{out}}\equiv(\ket{\Theta_{1}}\bra{\Theta_{1}}_{u}\otimes\mathbb{I}_{d})\circ\Pi_{2}^{L}\,,$ (46)
@@
Having fixed the kinematics, and the final state depends only on our choice for the initial flavor eigenstates, namely |$\ket{ij}_{ud}\in H_{f}$. We can thus identify the matrix elements of $\mathcal{S}_{f}$ from Sec. II.2:
@@ Appendix C Electroweak Scattering Amplitudes
We only consider incoming and outcoming fermionic states that are described by negative helicity spinors $u_{L}$ for particles and positive helicity spinors $v_{R}$ for anti-particles. In the Dirac representation, the four-component spinors in the helicity basis are: [^10] The spinors are eigenstates of the helicity operator $\hat{\lambda}=\frac{\vec{\Sigma}\cdot\vec{p}}{|\vec{p}|}\,,\qquad\Sigma_{i}=\frac{1}{2}\begin{pmatrix}\sigma_{i}&0\\ 0&\sigma_{i}\end{pmatrix}\,,$ where $\vec{\Sigma}$ is the spin operator and $\sigma_{i}$ are the Pauli matrices.
@@
where the different signs between the charged- and neutral-currents comes from fermion crossing. For the process obtained by crossing symmetry
@@
Throughout the analysis we employed for the QED and EW couplings the corresponding 1-loop RG evolution. The numerical boundary conditions are taken at $\sqrt{s}=m_{Z}=91.187~{}\rm{GeV}$ to be $\alpha_{1}=0.0101$ and $\alpha_{2}=0.034$.

### v2

@@ Appendix A Description of Scattering/Measurement Process
The $\mathcal{S}$ operator is unitary and it can be defined in terms of the transfer operator $\mathcal{T}$:
@@
    $\Pi_{\rm{out}}$ corresponds to the final state measurement. First, we restrict the final state to have two particles [14] with negative helicity, a combination of actions denoted by $\Pi_{2}^{L}$. Then, we measure the scattering angle $\Theta$ of one of the final state particles, which without loss of generality we choose to be the up-type quark. [^11] Even though the scattering amplitude is independent of the azimuthal angle $\Phi$, an actual measurement of the final state particle in real space would involve the placement of a detector at a definite $\Phi$ value. We assume that such an angle is arbitrarily selected in our idealized measurement procedure. The full measurement can be written schematically as:
    $\Pi_{\rm{out}}\equiv(\ket{\Theta}\bra{\Theta}_{u}\otimes\mathbb{I}_{d})\circ\Pi_{2}^{L}\,,$ (46)
@@
Having fixed the kinematics, the final state depends only on our choice for the initial flavor eigenstates, namely |$\ket{ij}_{ud}\in H_{f}$. We can thus identify the matrix elements of $\mathcal{S}_{f}$ from Sec. II.2:
@@ Appendix C Electroweak Scattering Amplitudes
We only consider incoming and outgoing fermionic states that are described by negative helicity spinors $u_{L}$ for particles and positive helicity spinors $v_{R}$ for anti-particles. In the Dirac representation, the four-component spinors in the helicity basis are: [^12] The spinors are eigenstates of the helicity operator $\hat{\lambda}=\frac{\vec{\Sigma}\cdot\vec{p}}{|\vec{p}|}\,,\qquad\Sigma_{i}=\frac{1}{2}\begin{pmatrix}\sigma_{i}&0\\ 0&\sigma_{i}\end{pmatrix}\,,$ where $\vec{\Sigma}$ is the spin operator and $\sigma_{i}$ are the Pauli matrices.
@@
where the different signs between the charged- and neutral-currents comes from fermion crossing. Note that since we are fixing the external fermion masses to their current values and we perform the bulk of our analysis below the top mass, contributions induced by Higgs exchange that are proportional to the small fermion masses are negligible and we, therefore, omit them.
For the process obtained by crossing symmetry
@@
where $u_{L}(q^{i};p_{I})$ denotes the spinor in Eq. (C) for the quark $q$ of flavor $i$ and momentum $p_{I}$.
@@
Throughout the analysis, we employ for the QED and EW couplings the corresponding one-loop RG evolution. The numerical boundary conditions are taken at $\sqrt{s}=m_{Z}=91.187~{}\rm{GeV}$ to be $\alpha_{1}=0.0101$ and $\alpha_{2}=0.034$.